Is there any lower limit to the electric force between two particles placed at a certain distance?

Grade:12th pass

1 Answers

Rituraj Tiwari
3 years ago
Yes, there is a lower limit to the electric force between 2 particles placed at a separation of 1 cm.
We know that force ( sayF) between two charged particles ( having chargesqandQ) separated by a distance ofris given by :
F=kQq/r2
To find the lower limit we have to assume the charge on the both of the particles lowest possible and the quantum of charge is still the charge on an electron (saye).
Implies, lower limit of force in our case is :
F=ke2/r2
Hereris1cm=0.01m
k=9×109N−m2/C2
Implies
F (Lowest possible)=2.3×10−24N